# Uniform convergence of the smooth calibration error and its relationship with functional gradient

### v2

@@ 1 Introduction
Probabilistic prediction plays a central role in many real-world applications, such as healthcare [31], weather forecasting [44], and language modeling [45], where uncertainty estimates are often critical. Since ensuring the reliability of such predictions has become a key challenge in modern machine learning, calibration [Dawid1982, foster1998asymptotic]—which requires that predicted probabilities align with the actual frequency of the true label—has attracted increasing attention. Calibration is a relatively weak condition that can be satisfied by simple models. However, it is frequently violated in practice. Notably, guo2017calibration demonstrated that many modern deep learning models can be significantly miscalibrated, which has since drawn increasing attention to calibration within the machine learning community. Although various regularization techniques [pmlr-v80-kumar18a, karandikar2021soft, NEURIPS2022_33d6e648, marx2023calibration] and recalibration methods [zadrozny2001obtaining, guo2017calibration, gupta21b, sun2023minimumrisk, NEURIPS2019_8ca01ea9] have been proposed to improve calibration, most of them are either empirically evaluated without theoretical guarantee or lead to a trade-off between calibration and sharpness [Kuleshov2015], degrading predictive accuracy. As a result, it remains unclear which learning algorithms can train well-calibrated predictors without sacrificing accuracy.
@@
To address this gap, we focus on the smooth calibration error (CE) [blasiok2023unify], a recently proposed calibration metric for binary classification that has favorable theoretical properties. Using smooth CE, we contribute to answering the above question through two main theoretical advances. First, we derive a uniform convergence bound for smooth CE, showing that the population-level smooth CE can be bounded by the smooth CE over the training dataset plus a generalization gap (Section 3). Second, we prove that such training smooth CE can be bounded by the norm of the functional gradient (or its approximation) of the loss evaluated on training data, providing a principled criterion for optimizing calibration (Section 4). Taken together, these results show that algorithms that achieve small functional gradients while regularizing model complexity can obtain small smooth CE.
We apply our theoretical framework to three representative algorithms for the binary classification closely tied to functional gradients: gradient boosting trees (GBTs), kernel boosting, and two-layer neural networks (NNs). For GBTs, we provide the theoretical guarantee for their calibration performance, showing that the smooth CE decreases with iteration (Section 4.1). For kernel boosting, we analyze the trade-off between optimization and model complexity induced by the reproducing kernel Hilbert space (Section 4.2). For two-layer NNs, we leverage their connection to the neural tangent kernel to study calibration behavior on the basis of function space optimization (Section 4.3). Furthermore, for kernel boosting and two-layer NNs, by introducing a margin-based assumption, we derive sufficient conditions on the sample size and the number of iterations required to simultaneously achieve $\epsilon$-level smooth CE and misclassification rate under appropriately chosen hyperparameters.
@@ 2.1 Problem setting of binary classification
[…]
We consider a binary classification problem in the supervised setting. Let $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$ denote the data domain, where $\mathcal{X}\subset{}^{d}$ is the input space and $\mathcal{Y}=\{0,1\}$ is the label space. We assume that data points are sampled from an unknown data-generating distribution $\mathcal{D}$ over $\mathcal{X}\times\mathcal{Y}$. Let $\mathcal{F}$ be a class of predictors $f:\mathcal{X}\to[0,1]$ approximating the ground-truth conditional probability $f^{*}(x)=\mathbb{E}[Y|x]$.
@@ 2.2 Calibration metrics and loss functions
Given a distribution $\mathcal{D}$ and a predictor $f$, we say the model is perfectly calibrated if $\mathbb{E}[Y\mid f(X)]=f(X)$ holds almost surely [blasiok2023unify]. Various metrics have been proposed to quantify the deviation from perfect calibration. The most widely used metric is the expected calibration error (ECE): $\mathrm{ECE}(f):=\mathbb{E}\left[\left|\mathbb{E}[Y\mid f(X)]-f(X)\right|\right].$ Despite its popularity, ECE is difficult to estimate efficiently [arrieta2022metrics, lee2023t] and is known to be discontinuous [foster2018smooth, kakade2008deterministic]. As a more tractable alternative, we focus on the smooth CE [blasiok2023unify]:
@@
Let $\mathrm{Lip}_{L}([0,1],[-1,1])$ be the set of $L$-Lipschitz functions from $[0,1]$ to $[-1,1]$. Given $f$ and $\mathcal{D}$, the smooth CE is defined as
@@
We also express $\mathrm{smCE}(f,\mathcal{D})=\sup_{h\in\mathrm{Lip}_{1}([0,1],[-1,1])}\left\langle h(f(X)),Y-f(X)\right\rangle_{L_{2}(\mathcal{D})}$. The smooth CE offers favorable properties such as continuity and computational tractability [blasiok2023unify, hu2024testing]. Moreover, the smooth CE provides both upper and lower bounds on the binning ECE, making it a principled surrogate for analyzing calibration. See Section 5 for details.
@@
where $\mathrm{Lip}_{1}(\real,[-4,4])$ defines a class of $1$-Lipschitz functions from to $[-4,4]$ and the choice of $4$ reflects the fact that the sigmoid function is $1/4$-Lipschitz. Then, this post-processing leads to defining a smooth CE directly over the logit function:
@@
Similarly to Eq. (3), the following relation holds:
@@ 3 Uniform convergence of the smooth CE
As shown in Eqs. (3) and (6), (dual) smooth CE is related to improvements in the population risk, and blasiok2023when utilized these relationships to analyze when algorithms can achieve a small smooth CE. However, their results are defined over population-level quantities, where the expectation by $\mathcal{D}$ is taken, making them inapplicable to practical algorithms trained on finite data points.
Since our goal is to theoretically analyze the behavior of the smooth CE under standard learning algorithms trained on finite data, it is natural to consider its empirical counterpart. Given a dataset $S_{n}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ consisting of independently and identically distributed (i.i.d.) data points sampled from the data distribution $\mathcal{D}$, we define the empirical smooth CE [blasiok2023unify] given $S_{n}$ as
@@
Therefore, we need to evaluate $|\mathrm{smCE}(f,S_{\mathrm{te}})-\mathrm{smCE}(f,S_{\mathrm{tr}})|$, which we refer to as the smooth CE generalization gap. Combining this gap with Eq. (8), we obtain the desired bound. To evaluate the smooth CE generalization gap, we use the covering number bound. Suppose that $\mathcal{F}$ is equipped with the metric $\|\cdot\|_{\infty}$. Let $\mathcal{N}(\epsilon,\mathcal{F},\|\cdot\|_{\infty})$ be an $\epsilon$-cover with metric $\|\cdot\|_{\infty}$ of $\mathcal{F}$, with the cardinality $N(\epsilon,\mathcal{F},\|\cdot\|_{\infty})$. Then, for any $f\in\mathcal{F}$, there exists $\tilde{f}\in\mathcal{N}(\epsilon,\mathcal{F},\|\cdot\|_{\infty})$ such that $\|f-\tilde{f}\|_{\infty}\leq\epsilon$.
@@
[…]
We provide a proof sketch to highlight the key differences from standard generalization bounds.
Proof sketch.
To address this, we introduce $\{\sigma_{i}\}$s in a way that preserves the structure of the linear convex formulation of smooth CE. Since the resulting bound does not take the form of standard Rademacher complexity, we discretize the hypothesis class $\mathcal{F}$ and derive a covering number bound. ∎
In certain cases, the Rademacher complexity offers a more interpretable characterization. The empirical and expected Rademacher complexities of a function class $\mathcal{F}$ are defined as ^$\hat{\mathfrak{R}}_{S}(\mathcal{F})\coloneqq\mathbb{E}_{\sigma}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}f(x_{i})$ , and $\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})\coloneqq\mathbb{E}_{\mathcal{D}^{n}}[\hat{\mathfrak{R}}_{S}(\mathcal{F})]$, respectively.
@@
$\sup_{f\in\mathcal{F}}|\mathrm{smCE}(f,S_{\mathrm{te}})-\mathrm{smCE}(f,S_{\mathrm{tr}})|\leq C_{1}\mathfrak{R}_{\mathcal{D},2n}(\mathcal{F})(\log 2n)^{2}+\frac{C_{2}}{\sqrt{n}}+\sqrt{\frac{\log\delta^{-1}}{n}}.$ (10)
Remark 1.
Since the smooth CE involves evaluating the composite function $h(f(X))$, standard chaining or Rademacher bounds would introduce complexity over the composite class $\mathrm{Lip}_{1}([0,1],[-1,1])\circ\mathcal{F}$, which is typically much richer than $\mathcal{F}$. Our analysis circumvents this dependence. Furthermore, unlike conventional bounds that derive covering number estimates from the Rademacher complexity, we proceed in the reverse direction, deriving a Rademacher-type bound from a covering number argument. This leads to an additional $(\log 2n)^{2}$ factor.
Combining Theorem 1 and Corollary 1 with Eq. (8), we obtain the desired bound.
@@
Under the same assumptions as in Theorem 1, there exist universal constants $\{C_{i}\}$s such that with probability at least $\!1-\delta$ over the draw of $S_{\mathrm{tr}}$, each of the following holds for all $f\!\in\!\mathcal{F}$:
[…]
$|\mathrm{smCE}(f,\mathcal{D})-\mathrm{smCE}(f,S_{\mathrm{tr}})|$ $\leq\frac{C_{4}}{\sqrt{n}}+C_{1}\mathfrak{R}_{\mathcal{D},2n}(\mathcal{F})(\log 2n)^{2}+\sqrt{\frac{2\log\frac{3}{\delta}}{n}}.$ (12)
Thus, controlling the complexity of the hypothesis class and reducing the training smooth CE guarantee a small population smooth CE. We next explore how to achieve a small training smooth CE using a functional gradient.
@@ 4 Controlling the smooth CE via functional gradient
In this section, we discuss how we can minimize the training smooth CE. From Eqs. (3) and (6), (dual) smooth CE can be characterized via the (dual) post-processing gap, which describes the potential improvement over functions. Building on this perspective, we show that both smooth and dual smooth CEs can be further characterized using functional gradients. Since we focus on the functional gradients over the training dataset, they are just finite-dimensional vectors. We provide the precise connection between functional gradients and the population smooth CE in Appendix G.
@@
The connection between functional gradients and the smooth CE can be extended to general proper scoring rules using the post-processing gaps, see Appendix G for the details.
@@
In this section, we express the empirical risk as $L_{n}(g)=\frac{1}{n}\sum_{i=1}^{n}\ell_{\mathrm{ent}}(\sigma(g(X_{i})),Y_{i})$ and the functional gradient over the training dataset as $\nabla_{g}L_{n}(g)=(\nabla_{g}\ell_{\mathrm{ent}}(\sigma(g(X_{1})),Y_{1})),\dots,\nabla_{g}\ell_{\mathrm{ent}}(\sigma(g(X_{n})),Y_{n}))\in{}^{n}$.
@@ 4.1 Case study I: Gradient Boosting Tree
Gradient boosting [Friedmangreedy2001] is a widely used technique in machine learning for constructing predictive models by iteratively adding base learners to minimize a given loss function. This method generates a sequence of functions $\{g^{(t)}\}_{t=0}^{T}$ through iterative updates of the form, for $t=0,\dots,T-1$:
$g^{(t+1)}(x)=g^{(t)}(x)+w_{t}h_{t}(x),$ (15)
where $w_{t}>0$ is a stepsize and $h_{t}$ is a function that approximates the negative functional gradient of the empirical loss [mason1999boosting]. If $h_{t}$ were chosen as the exact functional gradient, the resulting model could easily overfit the training data. Therefore, in practice, $h_{t}$ is often restricted to lie in a predefined function class, providing implicit regularization. Although gradient boosting methods are often empirically observed to be well-calibrated [niculescu2005obtaining, wenger2020non], their theoretical validation remains less understood. Our theoretical framework naturally applies to their analysis.
We approximate the functional gradient over the training dataset using a base learner $\{\psi_{\theta}\}$:
$\theta_{t}=\operatorname*{\mathop{\rm argmin}}_{\theta\in\Theta}\langle\nabla_{g}L_{n}(g^{(t)}),\psi_{\theta}\rangle_{L_{2}(S_{n})},$ (16)
where $\left\langle\nabla_{g}L_{n}(g^{(t)}),\psi_{\theta}\right\rangle_{L_{2}(S_{n})}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\nabla_{g}\ell_{\mathrm{ent}}(\sigma(g(X_{i})),Y_{i}))\psi_{\theta}(X_{i})$. In this work, we focus on $L_{1}$ gradient boosting [Friedmanlogistic, Peter2007, Zhang_2023], which adopts the following model update:
$g^{(t+1)}(x)=g^{(t)}(x)+w_{t}(\psi_{\theta_{t}}(x)-g^{(t)}(x)).$ (17)
This formulation represents a convex combination of base learners $\{\psi_{\theta}\}$. The complete procedure is outlined in Algorithm 1 in Appendix D, which also contains the formal proofs of this section.
Here, we focus on the gradient boosting tree [friedman2002stochastic, hastie2005elements], where we use a binary regression tree for $\{\psi_{\theta}\}$. A binary regression tree of depth $m$ partitions the input space $\mathbb{R}^{d}$ into at most $J\leq 2^{m}$ disjoint regions: $\mathbb{R}^{d}=\bigcup_{j=1}^{J}R_{j},\quad\text{with }R_{j}\cap R_{k}=\emptyset\text{ for }j\neq k.$ Each region $R_{j}$ is assigned a constant value $c_{j}\in\mathbb{R}$, defining a piecewise constant function: $\psi_{\theta}(x)=\sum_{j=1}^{J}c_{j}\cdot\mathbbm{1}_{\{x\in R_{j}\}},$ where $\theta=\{c_{j},R_{j}\}_{j=1}^{J}$.
Under this setting, the following is our main result for the training dual smooth CE:
@@
Given $S_{\mathrm{tr}}$, assume that $g^{(0)}=0$, $T\geq 2$, $w_{t}=1/(t+2)$, and $\sup_{\theta}\|\psi_{\theta}\|_{\infty}\leq B$ with $B\geq 2$. When $m\geq d$, there exists a universal constant $C_{5}$ such that
$\min_{t\in\{1,\dots,T\}}\mathrm{smCE}^{\sigma}(g^{(t)},S_{\mathrm{tr}})\leq\frac{C_{5}B^{2}}{T+2}.$ (18)
This result implies that using a binary tree with sufficient depth leads to a small training smooth CE. A binary tree with depth $m\geq d$ can partition $\mathbb{R}^{d}$ into hyperrectangles, allowing it to represent indicator functions over $\mathbb{R}^{d}$. Consequently, it can approximate the Lipschitz functions used in defining the dual smooth CE in Eq. (14). This assumption is therefore essential for effectively controlling the smooth CE. For an alternative assumption that allows us to bound the smooth CE, see Section 4.2.
Next, we provide a guarantee for the population smooth CE.

### v4

@@ 1 Introduction
Probabilistic prediction plays a central role in many real-world applications, such as healthcare [27], weather forecasting [42], and language modeling [43], where uncertainty estimates are often critical. Since ensuring the reliability of such predictions has become a key challenge in modern machine learning, calibration [Dawid1982, foster1998asymptotic]—which requires that predicted probabilities align with the actual frequency of the true label—has attracted increasing attention. Calibration is a relatively weak condition that can be satisfied by simple models. However, it is frequently violated in practice. Notably, guo2017calibration demonstrated that many modern deep learning models can be significantly miscalibrated, which has since drawn increasing attention to calibration within the machine learning community. Although various regularization techniques [pmlr-v80-kumar18a, karandikar2021soft, NEURIPS2022_33d6e648, marx2023calibration] and recalibration methods [zadrozny2001obtaining, guo2017calibration, gupta21b, NEURIPS2019_8ca01ea9] have been proposed to improve calibration, most of them are either empirically evaluated without a theoretical guarantee or lead to a trade-off between calibration and sharpness [Kuleshov2015], degrading predictive accuracy. As a result, it remains unclear which learning algorithms can train well-calibrated predictors without sacrificing accuracy.
@@
To address this gap, we focus on the smooth calibration error (CE) [blasiok2023unify, foster2018smooth, kakade2008deterministic], a recently proposed calibration metric for binary classification that has favorable theoretical properties. Using smooth CE, we contribute to answering the above question through two main theoretical advances. First, we derive a uniform convergence bound for smooth CE, showing that the population-level smooth CE can be bounded by the smooth CE over the training dataset plus a generalization gap (Section 3). Second, we prove that such training smooth CE can be bounded by the norm of the functional gradient (or its approximation) of the loss evaluated on training data, providing a principled criterion for optimizing calibration (Section 4). Taken together, these results show that algorithms that achieve small functional gradients while regularizing model complexity can obtain small smooth CE.
We apply our theoretical framework to three representative algorithms for the binary classification closely tied to functional gradients: gradient boosting trees (GBTs), kernel boosting, and two-layer neural networks (NNs). For GBTs, we provide the theoretical guarantee for their calibration performance, showing that the smooth CE decreases with iteration (Section 4.1). For kernel boosting, we analyze the trade-off between optimization and model complexity induced by the reproducing kernel Hilbert space (Section 4.2). For two-layer NNs, we leverage their connection to the neural tangent kernel to study calibration behavior on the basis of function space optimization (Section 4.3). Furthermore, for all these algorithms, by introducing a margin-based assumption, we derive sufficient conditions on the sample size and the number of iterations required to simultaneously achieve $\epsilon$-level smooth CE and misclassification rate under appropriately chosen hyperparameters.
@@ 2.1 Problem setting of binary classification
We denote random variables by capital letters (e.g., $X$) and deterministic values by lowercase letters (e.g., $x$). The Euclidean inner product and norm are denoted by $\cdot$ and $\|\cdot\|$, respectively.
We consider a binary classification problem in the supervised setting. Let $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$ denote the data domain, where $\mathcal{X}\subset\real^{d}$ is the input space and $\mathcal{Y}=\{0,1\}$ is the label space. We assume that data points are sampled from an unknown data-generating distribution $\mathcal{D}$ over $\mathcal{X}\times\mathcal{Y}$. Let $\mathcal{F}$ be a class of predictors $f:\mathcal{X}\to[0,1]$ approximating the ground-truth conditional probability $f^{*}(x)=\mathbb{E}[Y|x]$.
@@ 2.2 Calibration metrics and loss functions
Given a distribution $\mathcal{D}$ and a predictor $f$, we say the model is perfectly calibrated if $\mathbb{E}[Y\mid f(X)]=f(X)$ holds almost surely [blasiok2023unify]. Various metrics have been proposed to quantify the deviation from perfect calibration. The most widely used metric is the expected calibration error (ECE): $\mathrm{ECE}(f):=\mathbb{E}\left[\left|\mathbb{E}[Y\mid f(X)]-f(X)\right|\right].$ Despite its popularity, ECE is difficult to estimate efficiently [arrieta2022metrics, lee2023t] and is known to be discontinuous [foster2018smooth, kakade2008deterministic]. As a more tractable alternative, we focus on the smooth CE:
@@
Let $\mathrm{Lip}_{L}([0,1],[-1,1])$ be the set of $L$-Lipschitz functions from $[0,1]$ to [1,1][-1,1]. Given $f$ and $\mathcal{D}$, the smooth CE is defined as
@@
We remark that the concept of the smooth CE is also introduced in foster2018smooth, kakade2008deterministic. We also express $\mathrm{smCE}(f,\mathcal{D})=\sup_{h\in\mathrm{Lip}_{1}([0,1],[-1,1])}\left\langle h(f(X)),Y-f(X)\right\rangle_{L_{2}(\mathcal{D})}$. The smooth CE offers favorable properties such as continuity and computational tractability [blasiok2023unify, hu2024testing]. Moreover, the smooth CE provides both upper and lower bounds on the binning ECE, making it a principled surrogate for analyzing calibration. See Section 5 for details.
@@
where $\mathrm{Lip}_{1}(\real,[-4,4])$ defines a class of $1$-Lipschitz functions from to [4,4][-4,4] and the choice of $4$ reflects the fact that the sigmoid function is $1/4$-Lipschitz. Then, this post-processing leads to defining a smooth CE directly over the logit function:
@@
Similarly to Eq. equation 3, the following relation holds:
@@ 3 Uniform convergence of the smooth CE
As shown in Eqs. equation 3 and equation 6, (dual) smooth CE is related to improvements in the population risk, and blasiok2023when utilized these relationships to analyze when algorithms can achieve a small smooth CE. However, their results are defined over population-level quantities, where the expectation by $\mathcal{D}$ is taken, making them inapplicable to practical algorithms trained on finite data points.
Since our goal is to analyze the behavior of the smooth CE under standard learning algorithms trained on finite data, it is natural to consider its empirical counterpart. Given a dataset $S_{n}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ consisting of independently and identically distributed (i.i.d.) data points sampled from the data distribution $\mathcal{D}$, we define the empirical smooth CE [blasiok2023unify] given $S_{n}$ as
@@
Therefore, we need to evaluate $|\mathrm{smCE}(f,S_{\mathrm{te}})-\mathrm{smCE}(f,S_{\mathrm{tr}})|$, which we refer to as the smooth CE generalization gap. Combining this gap with Eq. equation 8, we obtain the desired bound. To evaluate the smooth CE generalization gap, we use the covering number bound. Suppose that $\mathcal{F}$ is equipped with the metric $\|\cdot\|_{\infty}$. Let $\mathcal{N}(\epsilon,\mathcal{F},\|\cdot\|_{\infty})$ be an $\epsilon$-cover with metric $\|\cdot\|_{\infty}$ of $\mathcal{F}$, with the cardinality $N(\epsilon,\mathcal{F},\|\cdot\|_{\infty})$. Then, for any $f\in\mathcal{F}$, there exists $\tilde{f}\in\mathcal{N}(\epsilon,\mathcal{F},\|\cdot\|_{\infty})$ such that $\|f-\tilde{f}\|_{\infty}\leq\epsilon$.
@@
$\!\sup_{f\in\mathcal{F}}|\mathrm{smCE}(f,S_{\mathrm{te}})\!-\!\mathrm{smCE}(f,S_{\mathrm{tr}})|\!\leq\!\inf_{\epsilon\geq 0}8\epsilon\!+\!24\!\int_{\epsilon^{\prime}}^{1}\!\sqrt{\frac{\ln N(\epsilon^{\prime},\mathcal{F},\|\cdot\|_{\infty})}{n}}\!d\epsilon^{\prime}\!+\!2\sqrt{\frac{\log\delta^{-1}}{n}}.$ (9)
Since the smooth CE involves evaluating the composite function $h(f(X))$, standard chaining bounds would introduce complexity over the composite class $\mathrm{Lip}_{1}([0,1],[-1,1])\circ\mathcal{F}$. However, by leveraging the smoothness property of smooth CE, the above bound does not include the complexity of Lipschitz functions. Combining Theorem 1 with Eq. equation 8, we obtain the bound for the population smooth CE.
@@
Under the same assumptions as in Theorem 1, there exist a universal constant $C_{1}$ such that with probability at least $1-\delta$ over the draw of $S_{\mathrm{tr}}$, the following holds for all $f\!\in\!\mathcal{F}$:
$|\mathrm{smCE}(f,\mathcal{D})-\mathrm{smCE}(f,S_{\mathrm{tr}})|$ $\leq\frac{C_{1}}{\sqrt{n}}+\inf_{\epsilon\geq 0}8\epsilon\!+\!24\!\int_{\epsilon^{\prime}}^{1}\!\sqrt{\frac{\ln N(\epsilon^{\prime},\mathcal{F},\|\cdot\|_{\infty})}{n}}d\epsilon^{\prime}\!+\!3\sqrt{\frac{\log\frac{3}{\delta}}{n}}.$ (10)
The first term of the right-hand side represents the complexity of the Lipschitz function class.
In certain cases, the Rademacher complexity offers a more interpretable characterization. The empirical and expected Rademacher complexities of a function class $\mathcal{F}$ are defined as $\hat{\mathfrak{R}}_{S}(\mathcal{F})\coloneqq\mathbb{E}_{\sigma}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}f(x_{i})$ , and $\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})\coloneqq\mathbb{E}_{\mathcal{D}^{n}}[\hat{\mathfrak{R}}_{S}(\mathcal{F})]$, respectively.
@@
Under the same assumptions as in Theorem 1, there exist a universal constant $C_{2}$ such that with probability at least $1-\delta$ over the draw of $S_{\mathrm{tr}}$, the following holds:
$\sup_{f\in\mathcal{F}}|\mathrm{smCE}(f,\mathcal{D})-\mathrm{smCE}(f,S_{\mathrm{tr}})|$ $\leq\frac{C_{2}}{\sqrt{n}}+4\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})+2\sqrt{\frac{\log\frac{2}{\delta}}{n}}.$ (11)
The first term on the right-hand side reflects the complexity of $\mathrm{Lip}_{1}([0,1],[-1,1])$, while the second corresponds to $\mathcal{F}$. A direct application of uniform convergence theory would yield a bound in terms of $\mathfrak{R}_{\mathcal{D},n}(\mathrm{Lip}_{1}([0,1],[-1,1])\circ\mathcal{F})$, which cannot be reduced to $\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})$ using the standard contraction lemma. Our analysis circumvents this difficulty by employing a covering argument for Lipschitz functions. Similarly, Rademacher complexity can be used to bound $|\mathrm{smCE}(f,S_{\mathrm{te}})-\mathrm{smCE}(f,S_{\mathrm{tr}})|$. However, such a bound depends on the complexity of the Lipschitz function, making it suboptimal relative to Theorem 1. See Appendix D for a detailed comparison between Corollary 1 and Theorem 2.
In conclusion, by jointly controlling the complexity of the hypothesis class and reducing the training smooth CE, we guarantee a small population smooth CE. In the next section, we show how to achieve a small training smooth CE via functional gradients.
@@ 4 Controlling the smooth CE via functional gradient
In this section, we discuss how we can minimize the training smooth CE. From Eqs. equation 3 and equation 6, (dual) smooth CE can be characterized via the (dual) post-processing gap, which describes the potential improvement over functions. Building on this perspective, we show that both smooth and dual smooth CEs can be further characterized using functional gradients. Since we focus on the functional gradients over the training dataset, they are just finite-dimensional vectors. We provide the precise connection between functional gradients and the population smooth CE in Appendix H.
@@
The connection between functional gradients and the smooth CE can be extended to general proper scoring rules using the post-processing gaps, see Appendix H for the details.
@@
In this section, we express the empirical risk as $L_{n}(g)=\frac{1}{n}\sum_{i=1}^{n}\ell_{\mathrm{ent}}(\sigma(g(X_{i})),Y_{i})$ and the functional gradient over the training dataset as $\nabla_{g}L_{n}(g)=(\nabla_{g}\ell_{\mathrm{ent}}(\sigma(g(X_{1})),Y_{1})),\dots,\nabla_{g}\ell_{\mathrm{ent}}(\sigma(g(X_{n})),Y_{n}))\in\real^{n}$.
@@ 4.1 Case study I: Gradient Boosting Tree
Gradient boosting [Friedmangreedy2001] is a widely used method for constructing predictive models by iteratively adding base learners to minimize a loss function. It generates a sequence of functions $\{g^{(t)}\}_{t=0}^{T}$ via the updates, for $t=0,\dots,T-1$:
$g^{(t+1)}(x)=g^{(t)}(x)-w_{t}\psi_{t}(x),$ (14)
where $w_{t}>0$ is the stepsize and $\psi_{t}$ approximates the functional gradient of the empirical loss [mason1999boosting]. Since using the exact functional gradient often leads to overfitting, $\psi_{t}$ is restricted to a predefined function class $\Psi$, which provides implicit regularization. We impose the following mild assumption on $\Psi$, which is readily satisfied by common choices such as regression trees.
Assumption 1.
The set of real-valued functions $\Psi$ satisfies: for every $\psi\in\Psi$, the negation $\psi$ also belongs to $\Psi$; $\sup_{\psi\in\Psi}\|\psi\|_{\infty}\leq B$ for some $B\geq 1$; and the constant function 0 is included in $\Psi$.
To obtain the update direction, we iteratively solve $\psi^{(t)}=\operatorname*{arg\,min}_{\psi\in\Psi}L_{n}\big(g^{(t)}-w_{t}\psi(x)\big)$. Many boosting methods approximate this by a quadratic upper bound, leading to the squared loss
$\psi_{t}=\operatorname*{arg\,min}_{\psi\in\Psi}\|Mw_{t}\psi-\nabla_{g}L_{n}(g^{(t)})\|^{2}_{L_{2}(S_{n})}=\frac{1}{n}\sum_{i=1}^{n}|Mw_{t}\psi(X_{i})-\nabla_{g}\ell_{\mathrm{ent}}(\sigma(g(X_{i})),Y_{i}))|^{2}.$ (15)
where $M$ is the smoothness parameter of the loss (e.g., $M=1/4$ for cross-entropy).
Although gradient boosting is often observed to be well calibrated in practice [niculescu2005obtaining, wenger2020non], its theoretical guarantees remain less understood. Our framework provides a natural way to analyze this. Following prior analysis on boosting, we impose a standard margin assumption:
Assumption 2 (telgarsky2013margins).
Let $\Delta_{n}=\{q\in\real_{+}^{n}\mid\sum_{i=1}^{n}q_{i}=1\}$ denote the $n$-dimensional probability simplex. Given a dataset $S_{\mathrm{tr}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$, there exists $\gamma>0$ such that for every $q\in\Delta_{n}$, there exists $\psi\in\Psi$ satisfying $\sum_{i=1}^{n}q_{i}(2Y_{i}-1)\psi(X_{i})\geq\gamma B$.
Here, $2Y_{i}-1$ maps the binary label $Y_{i}\in\{0,1\}$ to $\pm 1$. This condition guarantees that for any weighted sample distribution $q\in\Delta_{n}$, one can choose a base learner with nontrivial correlation to the labels. Such margin assumptions are also standard in classification theory [wei2018margin]. Under this setup, we present our main result for the training dual smooth CE:
@@
Given $S_{\mathrm{tr}}$, under Assumptions 1 and 2, if $T\geq 2$ with constant stepsize $w_{t}=w$, the averaged predictor $\bar{g}^{(T)}=\frac{1}{T}\sum_{t=0}^{T-1}g^{(t)}$ satisfies
$\mathrm{smCE}^{\sigma}\big(\bar{g}^{(T)},S_{n}\big)\leq\frac{L_{n}(g^{(0)})}{\gamma BwT}+\frac{wB}{8\gamma}.$ (16)
Proof outline.
By definition, the smooth CE is bounded by the $L_{1}$ norm of the functional gradient:
$\mathrm{smCE}^{\sigma}(g,S_{\mathrm{tr}})\leq\frac{1}{n}\sum_{i=1}^{n}|\nabla_{g}\ell_{\mathrm{ent}}(g(X_{i}),Y_{i})|=\|\nabla_{g}\ell_{\mathrm{ent}}(g(X),Y)\|_{L_{1}(S_{n})}.$ (17)
Under Assumption 2, this norm can be further bounded as
$\|\nabla_{g}\ell_{\mathrm{ent}}(g(X),Y)\|_{L_{1}(S_{n})}\leq\frac{1}{\gamma B}\big\langle\psi_{t},\nabla_{t}\big\rangle_{L_{2}(S_{n})}+\frac{w}{8\gamma}B$ (18)
as shown in Lemma 4 in Appendix E. The inner product $\big\langle\psi_{t},\nabla_{t}\big\rangle_{L_{2}(S_{n})}$ is then bounded based on the standard convex optimization techniques. See Appendix E.3 for the complete proof. ∎
Thus, choosing $w=\mathcal{O}(1/\sqrt{T})$ ensures that the training smooth CE converges to 0 at rate $\mathcal{O}(1/\sqrt{T})$.
Next, we establish a guarantee for the population smooth CE by applying Theorem 2. This requires specifying $\Psi$ to evaluate the complexity of $\mathcal{F}$. We consider gradient boosting trees [friedman2002stochastic, hastie2005elements], where $\Psi$ is the class of binary regression trees. A binary regression tree of depth $m$ partitions the input space $\mathbb{R}^{d}$ into at most $J\leq 2^{m}$ disjoint regions: $\mathbb{R}^{d}=\bigcup_{j=1}^{J}R_{j},\text{with }R_{j}\cap R_{k}=\emptyset\text{ for }j\neq k.$ Each region $R_{j}$ is assigned a constant $c_{j}\in\mathbb{R}$, yielding a piecewise constant function: $\psi_{\theta}(x)=\sum_{j=1}^{J}c_{j}\cdot\mathbbm{1}_{\{x\in R_{j}\}}$ with parameters $\theta=\{c_{j},R_{j}\}_{j=1}^{J}$. The complete GBT algorithm is provided in Appendix E. Under these settings, we obtain the following result: